A scuba tank is filled with air to a gauge pressure of 204 atm when the air temperature is 33 ∘C. A diver then jumps into the ocean and, after a short time on the ocean surface, checks the tank's gauge pressure and finds that it is only 196 atm. Assuming the diver has inhaled a negligible amount of air from the tank, what is the temperature of the ocean water? Express your answer to two significant figures and include the appropriate units.
Answer #1

given

P1 = 204 atm ,

T1 = 33oC

= 33 + 273

= 306 K

P2 = 196 atm ,

T2 = ?

we have equation and that the relation is

P1 / P2 = T1 / T2

T2 = T1 P2 / P1

T2 = 306 X 196 / 204

T2 = 294 K

T2 = 21oC

the temperature of the ocean water is T2 = 21oC